A rectangular swimming pool is surrounded by a walk. The area of the pool is 323 sq ft, and the outside dimensions of the walk are 20x22 ft. How wide is the walk?
I know this should be simple, but I'm struggling with it!
Thank you so much for your help, what ever kind person answers my question!
Lauren

Answer by scott8148(6628)

the walk is on all four sides of the pool, so the dimensions of the pool are 20-2w and 22-2w
__ where w is the width of the walk

(20-2w)(22-2w)=323 __ 440-84w+4w^2=323 __ 4w^2-84w+117=0

use quadratic formula to find w




answer below


w=1.5